47. Describe the six assumptions behind the economic order quantity model for stock control. Within ur answer, ensure to support each point by applying the assumption to a business example
Answer #1

Let's first understand the economic order quantity model:

Q = \sqrt{\frac{2 \times D\times S}{H}}

where symbols have the following meaning:

Q = optimal order quantity

D = Annual demand (nos. of units)

S = Cost to place a single order or setup cost

H = Carrying cost per unit

Sl. No. Assumption Support by example
1.

The cost to place an order or the set up cost (S) remains constant. The inventory carrying cost (H) is constant.

In a business scenario, the cost to place an order typically comprises of administrative costs of preparing purchase orders, service orders, contacting suppliers, evaluating suppliers, follow up phone calls, transportation costs, inspection, quality tests, process control, loading and unloading costs. Most of these costs are lumpsum costs. They may vary from order to order in real life. Cost towards preparing purchase orders, service orders, contacting suppliers, evaluating suppliers may be one time costs incurred just prior to the first order. In the subsequent order these costs may not be there. In that case, such costs should be averaged out across orders and the average cost should be used in the model.

The fluctuation in interest rate or warehousing cost may not keep the inventory carrying cost constant for the compnay.

2. The total demand (D) for the year is known, the demand per unit of time is known and demand is uniform throughout the year Any company involved in fast moving consumer goods production and sale will be most suitable example for such kind of assumption. The demands for items like soap, hair colour, shampoo don't see any seasonal or time period based fluctuations. Demands are pretty constant during the year and for the year. The is no major fluctuations. Hence, the demand rate for all the raw materials can remain constant throughout the year. Businesses with seasonal variations or fluctuations due to any other factors, may not have a constant demand rate throughout the year.
3. The lead time to process the order and complete the order is constant Under a long term contracts with suppliers, businesses can typically have a constant lead time. Unplanned disruptions like strike of transporters, non availability of labor force for loading - unloading etc may lead to longer lead time for a company.
4. The ordered quantity (Q) is received in time and in full quantity (i.e. in whole lot)

This is also a reasonable assumption to make however in following examples this may get violated:

1. When the quantity demanded is not able to complete full container load or full truck load. So, the part load may be delivered in second lot say after a few days when the supplier is able to aggregate orders from other customers to complete one full container or truck load.

5. No cash discounts, no settlement discounts The EOQ model assumes here is no discount in the purchase price. There are many businesses where there usually is the concept of volume discount. Once the quantity ordered exceeds a threshold value, the supplier offers some discount on the sale price of the subsequent orders. This is a fairly common practise in almost every business.
6. The optimal quantity for each raw material is determined separately Different raw materials will have different costs, demand patterns and consumptions. Hence, EOQ for each one of them may be different and hence simultaneous orders may not be possible. In real life this may not be the case. And hence, many a times, firms places orders for multiple raw materials simultaneously.